# Digital Durable Goods Monopoly

### v8

@@ Abstract
This paper develops a model of dynamic monopoly for digital durable goods characterized by two features: free disposability on the buyer side and zero marginal cost on the seller side. We show that when parties are sufficiently patient, the set of equilibrium seller payoffs contains the interval from the lowest-type buyer’s efficient surplus to the static monopoly commitment payoff, subject to the constraint that the lowest-type buyer receives an efficient allocation with probability one. The reason is that when allocative efficiency is set-valued rather than single-valued, Coasian forces do not generally pin down continuation values and therefore do not select a unique payoff outcome. These results show how structural properties of digital goods generate market indeterminacy.
@@
The high profitability of business-to-business digital goods markets makes them an important object of study [^1] See evidence in Appendix C.1. Business-to-business (B2B) digital goods, such as enterprise software, curated datasets, and access to APIs or AI models, are rarely purchased as standardized off-the-shelf products. Instead, they are acquired through negotiated contracts that bundle multiple attributes, including capacity, features, tiers, service levels, and price, into a single offer that the buyer either accepts or rejects. The bargaining process typically concerns custom pricing and licensing configurations, in which the vendor maintains a unified codebase while offering distinct modules or tiers of functionality. For example, a client in a regulated industry may require an “Advanced Compliance” module, whereas another client may view such features as unnecessary.
Two contractual features are salient in many digital B2B markets. First, buyers can underutilize purchased capacity or functionality and leave unused portions idle at essentially no cost, and realized usage is often non-contractible (Corrao et al., 2023). For example, a firm purchasing a dataset from a data broker such as Acxiom may query only the subset of consumer profiles relevant to a campaign. A corporation licensing enterprise software from Salesforce may leave entire modules (e.g., “Salesforce Maps”) idle. Second, once a digital product has been developed, expanding contractual access, such as enabling modules, granting seats, or raising usage limits, is close to costless at the margin relative to the fixed development cost (Quah, 2003; Goldfarb and Tucker, 2019). For data brokers such as Acxiom or Experian, after consumer records are aggregated and refined, providing filtered subsets or curated segments to clients involves minimal cost. Likewise, after firms such as Microsoft or OpenAI develop their core platforms, extending access to different feature sets, tiers, or API limits requires no significant resource expenditure. These features induce a distinct strategic structure. The buyer can accept an allocation that exceeds the efficient utilization level and discard the excess, while the seller can increase contracted access without incurring incremental production costs along the contractual allocation dimension. As a result, allocative efficiency is non-unique in a way that does not arise in standard physical-goods environments.
These negotiations also lie at the center of a classic theoretical problem: a monopolist with limited commitment selling a durable good to a buyer with private information. In canonical durable-goods monopoly models, this lack of commitment disciplines market power and can drive outcomes toward efficiency in the patient limit, as formalized by the Coase conjecture (Coase, 1972; Fudenberg et al., 1985; Gul et al., 1986). This paper uses a digitally motivated contracting environment to identify a boundary of this logic. When allocative efficiency is set-valued rather than single-valued, Coasian forces continue to impose constraints, but they do not generally pin down continuation values and hence do not select a unique payoff outcome.
Throughout the paper, “efficiency” refers to allocative efficiency in the allocation or quality dimension, conditional on trade, rather than efficiency in the timing dimension. [^2] With quality choice, equilibrium delay patterns are difficult to characterize in general (See an indeterminacy result in Appendix B.4). Accordingly, the main results focus on the set of seller payoffs sustainable in equilibrium and how this set changes when allocative efficiency is non-unique. The central question is therefore the following. What constraints does Coasian discipline impose when the set of efficient allocations is large, and which payoff outcomes can be sustained?
We formalize these questions in a dynamic monopoly model with one-sided private information. The buyer has a privately known type that captures the marginal benefit of utilization, or equivalently the negative of a marginal implementation cost. In each period, the seller makes a single take-it-or-leave-it offer consisting of an allocation–price pair, and the buyer either accepts or rejects. The allocation is interpreted as a contractual allocation, such as access rights, enabled features, seats, capacity, or precision, while realized usage is chosen by the buyer after purchase. Formally, the buyer may discard unused portions at zero cost (free disposability). On the seller side, we assume zero marginal cost along the contractual allocation dimension, so the seller’s profit equals the transfer. The single-offer protocol isolates intertemporal screening. It provides a reduced form for many B2B private-offer negotiations, in which the contracting object is a bundled proposal rather than a posted menu; Appendix C.2 discusses institutional support. Allowing within-period menus would reintroduce intratemporal screening and reduce the dynamic problem to a static menu benchmark, which we analyze separately in Appendix B.3.
@@
The main result is a Folk-Theorem-type attainability statement for seller payoffs. Let $u_{\text{min}}$ denote the surplus of the lowest-type buyer under the efficient allocation. Let $\pi^{e}(F)$ denote the seller’s optimal static payoff subject to the additional constraint that the lowest-type buyer receives an allocatively efficient allocation with probability one. We show that for every $\epsilon>0$, there exists $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$, the set of equilibrium seller payoffs contains the interval
$u_{\text{min}},\ \pi^{e}(F)-\epsilon].$
Moreover, these high-payoff equilibria can be constructed to satisfy an on-path Markov restriction on buyer behavior, a refinement that often collapses multiplicity in canonical Coasian settings. In particular, we do not claim that $\pi^{e}(F)$ is a universal upper bound across all equilibria in the general quality-choice environment. Appendix B.1 provides upper-bound results for broad classes of equilibria.
Free disposability and zero marginal contractual cost imply a sharp structural property. Each type has a unique efficient utilization level, but any contractual allocation above is allocatively efficient for that type because the buyer can discard slack. Consequently, the seller can change the contracted allocation path without reducing allocative efficiency. This decouples two roles that coincide in canonical Coasian models: (i) delivering efficiency and (ii) pinning down continuation values. When efficiency is essentially single-valued, it tightly restricts feasible continuation paths and often yields sharp predictions. When allocative efficiency is set-valued, the seller can select allocation paths that generate different continuation values while respecting the incentive constraints on the marginal type. This flexibility sustains the interval of attainable payoffs.
The role of “zero marginal cost” (together with free disposability) should be understood as a structural boundary rather than an innocuous approximation. The payoff indeterminacy disappears when allocative efficiency becomes single-valued, e.g., without free disposability or with strictly increasing marginal costs. Section 5 provides detailed results in discrete-type environments.
Thus, in digital contracting environments, the negotiated outcomes may be more sensitive to norms and bargaining practices, and less tightly pinned down by fundamentals than in canonical models.
Our set-valued efficiency perspective also helps clarify why Ausubel and Deneckere (1989b) obtain a folk theorem result, as a case in which the lowest type’s efficient set is non-singleton. A detailed discussion is deferred to Section 4.4.
@@
Section 2 presents the model. Section 3 characterizes static benchmarks. Section 4 establishes the Coasian equilibrium and the Folk-Theorem-type attainability result. Section 5 studies cases with unique efficient allocations. Section 6 discusses applications. A detailed literature review is deferred to Section 7. Section 8 concludes. The appendices provide proofs, additional results, and institutional discussion.
@@ 2.1 Primitives
The buyer is characterized by a privately known type $\theta\in\Theta=[\underline{\theta},\overline{\theta}]$, drawn from a cumulative distribution function $F(\cdot)$. [^3] In the baseline model, buyer types are drawn from a continuum to simplify the static commitment benchmark. We analyze finite-type distributions in Section 5 and Appendix B. The type $\theta$ may take negative values; for instance, it can be interpreted as c-c, representing marginal implementation or utilization costs.
A single good is available for trade. The allocation space is $X=[\underline{x},\overline{x}]$, where $\underline{x}>0$ denotes the minimal technologically feasible allocation. The variable $x$ may represent either the quantity or the quality of the good. We interpret $x$ as a contractual allocation (access/rights/capacity), while realized usage $x^{\prime}$ is chosen by the buyer and is not directly contractible.
@@
This formulation implies that although the buyer purchases an allocation $x$, she optimally selects an actual consumption level $x^{\prime}\in[0,x]$. Any excess allocation $x-x^{\prime}$ is discarded without cost. Under zero marginal cost along the contractual allocation dimension, the seller’s profit equals $p$.
@@
Assumption 3 requires that the virtual surplus at the efficient consumption level be nonnegative for every type $\theta$. This assumption is imposed for equilibrium existence in the continuum‑type, continuous‑action baseline and does not drive the results; Appendix B shows the main constructions in a discretized action space without Assumption 3.
@@ 2.2 Timing and Solution Concept
Acceptance terminates the game [^4] This is consistent with a durable access right (or a negotiated contract whose terms remain fixed over the relevant horizon).. If the buyer accepts in period $t$ ($a_{t}=1$), the buyer’s payoff is
@@
This constrained benchmark matches the efficiency requirement that emerges in our equilibrium constructions for the lowest type. Adapting the analysis from Lemma 3 yields the following characterization.
@@ 4.1 The Coasian Equilibrium
This outcome mirrors the standard Coasian benchmark in the following sense: every buyer type receives an efficient allocation with no delay, and the seller’s profit is driven to the surplus of the lowest buyer type. Since $u_{\min}<\pi(F)$, this payoff lies strictly below the static commitment benchmark.
@@ 4.3 The Folk Theorem
If $\underline{x}\geq x^{e}(\underline{\theta})$, then $\pi^{e}(F)=\pi(F)$ by Corollary 2. We therefore obtain the following corollary.
@@
Suppose $\underline{x}\geq x^{e}(\underline{\theta})$. For any discount factor $\delta$,
$u_{\min},\ \pi^{e}(F)]\ \supseteq\ SE(\delta).$
We discuss whether $\pi^{e}(F)$ constitutes the upper bound on all equilibrium payoffs in more general settings in Appendix B.
@@ 4.4 Comparison
In the framework of Ausubel and Deneckere (1989b), where the lowest buyer valuation is normalized to zero (the “no-gap” case), the equilibrium payoff set extends to the full commitment payoff. Our analysis provides a complementary perspective on this result. In the setting of Ausubel and Deneckere (1989b), the only allocation dimension is the discounted probability of trade. When the lowest type has valuation zero in the no-gap case, the efficient timing of trade is set-valued for this type, including infinite delay. Thus, under limited commitment, once the lowest type is allocated infinite delay, the allocations for other types may be arbitrary, provided they are monotone in type. In this context, the upper bound of the Folk Theorem can be interpreted as the maximum profit achievable subject to the constraint that the lowest-type buyer receives an efficient time of trade. In their setting, this constrained maximum coincides with the unconstrained static commitment payoff.
Robustness to Markov Refinements
We show that the Folk Theorem is robust to a class of Markov refinements that typically eliminate reputational equilibria and restore the Coase conjecture in standard settings. This distinguishes our environment from those benchmarks. We begin by recalling the definition of weak-Markov equilibria.
@@
We now introduce a relaxation of this concept.
@@
A common structural feature of these environments is that allocations are fixed, leaving price as the only screening instrument. In such settings, the on-path Markov refinement eliminates the seller’s ability to delay because the incentive to preempt future sales remains strong. In contrast, our framework allows the seller to choose both price and allocation, and the equilibrium path features decreasing allocations. This additional margin alters the incentives: accelerating sales requires offering the lower future allocation immediately, which reduces total surplus and therefore the seller’s profit. Hence, when agents are sufficiently patient, the incentive to maintain the equilibrium allocation sequence outweighs any gain from acceleration.
hen efficient allocations are unique.
@@
In this section, we analyze the case in which efficient allocations are unique. We adopt a discrete-type framework to show that both free disposability and zero marginal cost are necessary for the payoff multiplicity highlighted by the Folk Theorem. As discussed above, the skimming property generally fails in the presence of quality differentiation, which makes a direct characterization of equilibrium strategies difficult. Addressing the problem in full generality is therefore beyond the scope of this paper.
@@
We first show that the market indeterminacy, that is, equilibrium multiplicity, in Theorem 2 relies on free disposability. In particular, when free disposability is absent, the equilibrium set collapses to a single point.
@@
In Proposition 1, the unique equilibrium payoff converges to $\pi^{e}(F)$ in the patient limit. This occurs because, in the two-type case, the efficient outcome coincides with the static monopoly outcome subject to the constraint that the low type receives the efficient allocation. In this setting, the low type $\theta_{2}$ receives the efficient allocation. Moreover, since $\theta_{1}$ is the high type, its allocation is not subject to downward distortion in static monopoly outcome, and it therefore also receives the efficient level. On the other hand, when free disposability is present, the equilibrium payoff set spans $u_{\text{min}},\pi^{e}(F)]$. Thus, we have shown that equilibrium multiplicity relies on free disposability. Furthermore, attaining $\pi^{e}(F)$ is a special case driven by the two-type structure. In Appendix B.5, we show that $\pi^{e}(F)$ cannot be achieved in general without free disposability.
5.1 Positive Marginal Cost
@@
Taken together, these results show that the Folk Theorem does not arise solely from quality differentiation. In this sense, they suggest that the particular indeterminacy we study is most naturally associated with digital-goods environments.
pplications
Our primary applications include digital solutions and markets for data.
Digital Solutions.
Consider the market for digital solutions, including Software-as-a-Service (SaaS) platforms (e.g., Salesforce, Microsoft), specialized APIs (e.g., Stripe, Twilio), and AI models (e.g., OpenAI). After development, these firms negotiate large-scale subscriptions with enterprise buyers. Here, $x$ represents the volume of user licenses or the set of enabled features. A buyer of type $\theta$ selects a utilization level $x^{\prime}\leq x$ to enhance productivity or support collaboration, incurring an internal implementation cost $\theta x^{\prime}$ related to training, integration, and administration. [^9] Integration and administrative costs often scale approximately linearly with utilization. The buyer’s utility is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right],$
where $v(\cdot)$ is strictly concave, capturing diminishing marginal returns. [^10] For example, in a CRM platform, core features provide high marginal value, while incremental tools yield lower marginal benefits.
In this setting, free disposability implies that unused licenses or features may remain idle without penalty. Zero marginal cost reflects that, once the software infrastructure is in place, providing additional licenses or feature access entails negligible marginal expense for the seller.
Consumer Data.
A second application is the market for consumer data, in which brokers such as Acxiom and Experian sell curated datasets to business clients. Consider a continuum of consumer profiles indexed by $z$ and ordered by strictly decreasing valuation, where each profile $z$ generates a gross value $v^{\prime}(z)$. [^11] Equivalently, $v^{\prime}(z)$ may be interpreted as a normalized valuation. An allocation $x$ corresponds to a marketing list containing all profiles with valuations above the threshold $v^{\prime}(x)$. Thus, the list includes profiles from the highest-value $v^{\prime}(0)$ down to the marginal profile $v^{\prime}(x)$. For a downstream firm facing a constant marginal utilization cost $c=-\theta$, the realized profit is
$u(x,\theta)=\int_{0}^{x}\left(v^{\prime}(z)-c\right)^{+}\,\mathrm{d}z.$
In this framework, datasets are organized as threshold lists, which makes $x$ a one-dimensional quality measure. This reflects the practical constraint that arbitrary sampling or recombination of individual profiles is often costly or infeasible for the seller.
Data-Driven Decision Making.
Consider a decision-maker (e.g., an investor) facing a binary choice under uncertainty regarding a state $\omega\in\{-1,1\}$, with a uniform prior. A buyer of type $\theta$ obtains payoff $\theta a\omega$ from choosing an action $a\in\{-1,1\}$ when the action matches the true state. In the absence of data, expected utility is normalized to zero. The seller offers datasets indexed by $x\in[0,1/2]$, representing the maximal available signal precision.
Possession of a dataset with precision $x$ allows the buyer to conduct a symmetric experiment that updates the prior to posteriors (1/2x, 1/2+x)($^{\prime},\,1/2+x^{\prime}$) or (1/2+x, 1/2x)($^{\prime},\,1/2-x^{\prime}$), where the realized precision $x^{\prime}\leq x$ is chosen by the buyer. Generating this signal entails a cost $C(x^{\prime})$, where $C(\cdot)$ is strictly convex. [^12] For general references on uniform posterior-separable costs, see Frankel and Kamenica (2019), Denti et al. (2022), Denti (2022), and Caplin et al. (2022).
The buyer’s optimization problem is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[2\theta x^{\prime}-C(x^{\prime})\right].$
Defining $v(x^{\prime})\equiv-C(x^{\prime})$ and rescaling the type parameter to absorb the coefficient yields
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right],$
where $v(\cdot)$ is strictly concave. In this setting, free disposability corresponds to the buyer’s option to conduct a coarser experiment than the dataset permits, and zero marginal cost reflects that, once data is aggregated, providing higher-precision datasets entails no additional production cost.

### v10

# Coasian Dynamics with Set-Valued Allocative Efficiency: Information Goods

@@ Abstract
This paper studies a dynamic monopoly with limited commitment in which allocative efficiency is set-valued. The environment is characterized by two structural features: free disposability on the buyer side and zero marginal cost on the seller side along the allocation margin. We show that when parties are sufficiently patient, the set of equilibrium seller payoffs contains an interval ranging from the lowest-type buyer’s efficient surplus to a constrained static monopoly benchmark, subject to the requirement that the lowest-type buyer receives an efficient allocation with probability one. When efficiency is set-valued rather than single-valued, Coasian discipline operates only through an allocative constraint on the lowest type and does not generally pin down a unique equilibrium path. We interpret the results through the lens of information goods, which offer a clean benchmark in which allocative efficiency is set-valued.
@@
A central insight of the Coase conjecture is that limited commitment disciplines monopoly power in durable-goods markets. When a seller cannot commit to future terms, competition with her future self constrains equilibrium outcomes and leads to sharp predictions in canonical environments: market efficiency in the patient limit (Coase, 1972; Fudenberg et al., 1985; Gul et al., 1986). A key ingredient behind this selection logic is that allocative efficiency is essentially single-valued: in canonical single-margin environments, each type has a unique efficient timing of trade, namely immediate trade without discounting.
This paper identifies a structural boundary of that logic. We study dynamic monopoly environments in which allocative efficiency is set-valued rather than single-valued. In such settings, limited commitment continues to impose an efficiency constraint on the lowest-type buyer, but it does not generally pin down a unique equilibrium path. As a result, Coasian discipline operates only partially, and a range of equilibrium outcomes can be sustained in the patient limit.
One natural source of set-valued efficiency arises in the sale of information goods. Consider a seller who offers an information good at different levels of scope or precision, while the buyer decides ex post how much of the purchased information to use. A canonical example is a dataset. After raw data are aggregated and cleaned, the seller may offer datasets of varying breadth (e.g., larger lists of entries) or precision (e.g., additional attributes or finer segmentation). Unused records or unused precision can be freely disposed [^1] See Bergemann et al. (2018) and Corrao et al. (2023) for related formulations., so the buyer can accept a dataset that exceeds her efficient utilization without reducing realized utility. Moreover, once the underlying database has been created, expanding delivered scope or precision has zero marginal cost [^2] See Goldfarb and Tucker (2019) for a comprehensive review. for the seller. These features imply a simple property: although each type has a unique efficient utilization level, the set of allocatively efficient levels is typically large. Information goods therefore provide an extension of the Coasian durable-goods benchmark, in which allocative efficiency is generically set-valued rather than single-valued.
We formalize these forces in a dynamic monopoly model with one-sided private information. In each period, the seller makes a single take-it-or-leave-it offer $(x_{t},p_{t})$, where $x_{t}$ indexes the allocation level and $p_{t}$ is the transfer, and the buyer either accepts or rejects. Upon acceptance, the buyer chooses realized utilization $x^{\prime}_{t}\leq x_{t}$ and discards the remainder at no cost. The seller has zero marginal cost along the allocation margin, so profit equals the transfer. We use the single-offer protocol to isolate intertemporal screening under limited commitment. Allowing within-period menus reinstates intratemporal screening and reduces the dynamic problem to a static menu benchmark, which we analyze separately.
Results.
Throughout, efficiency refers to allocative efficiency in the allocation dimension, conditional on trade. We define a constrained static benchmark $\pi^{e}(F)$ as the maximal static payoff among incentive-compatible mechanisms that assign the lowest-type buyer an allocatively efficient allocation with probability one.
Our first result constructs a Coasian equilibrium that yields the seller the efficient surplus of the lowest-type buyer, denoted $u_{\min}$. This equilibrium provides a low-payoff continuation that can support higher-payoff outcomes.
@@ Results.
Our main result is a Folk-Theorem-type attainability statement for seller payoffs. For every $\epsilon>0$, there exists a discount factor $\underline{\delta}$ such that for all $\delta\geq\underline{\delta}$, the set of equilibrium seller payoffs contains the interval
$u_{\min},\ \pi^{e}(F)-\epsilon].$
Thus, as agents become patient, a continuum of seller payoffs is sustainable in equilibrium. Moreover, the high-payoff equilibria can be constructed to satisfy an on-path Markov restriction on buyer strategy—a refinement that often collapses equilibrium multiplicity in canonical Coasian environments.
The key force is structural. Free disposability and zero marginal cost along the allocation margin imply that allocative efficiency does not uniquely pin down the allocation level. Any allocation exceeding a type’s efficient utilization level is allocatively efficient because the buyer can freely dispose the slack and the seller incurs zero marginal cost. This non-uniqueness decouples two roles that coincide in canonical Coasian models: delivering efficiency and pinning down continuation equilibrium paths. As a result, the seller can vary future allocation levels while maintaining allocative efficiency for the buyer, which generates multiple self-enforcing continuation values. High-payoff screening paths can therefore be sustained by the credible threat of reverting to a low-payoff equilibrium continuation, such as the Coasian equilibrium.
One implication is that, in environments of the kind we model, equilibrium outcomes need not be tightly determined by primitives alone. Instead, equilibrium selection can depend on auxiliary features that coordinate continuation play (e.g., conventions or reputational considerations), even when underlying preferences and technology are unchanged.
We also provide results showing that this payoff indeterminacy disappears once allocative efficiency becomes single-valued, for example when buyers must consume the entire bought amount or when the seller faces strictly increasing marginal costs along the allocation margin. These results clarify that the key force in our setting is not quality differentiation per se, but the set-valued nature of allocative efficiency. Thus, this form of equilibrium indeterminacy is less likely to arise for physical goods: for a physical good (e.g., an automobile), delivering higher quality typically requires higher marginal cost, so allocative efficiency is single-valued along the quality margin.
Finally, the set-valued efficiency perspective offers a complementary lens through which to interpret classic Folk Theorem of the “no-gap” case.
@@
Section 2 presents the model. Section 3 characterizes static benchmarks. Section 4 establishes the Coasian equilibrium and the Folk-Theorem-type attainability result. Section 5 discusses examples, cases with unique efficient allocations, and menu offer per period. A detailed literature review is deferred to Section 6. Section 7 concludes. The appendices provide proofs and additional results.
@@ 2.1 Primitives
The buyer is characterized by a privately known type $\theta\in\Theta=[\underline{\theta},\overline{\theta}]$, drawn from a cumulative distribution function $F(\cdot)$. [^3] In the baseline model, buyer types are drawn from a continuum to simplify the static commitment benchmark. We analyze finite-type distributions in Section 5. The type $\theta$ may take negative values; for instance, it can be interpreted as c-c, representing marginal implementation or utilization costs.
A single good is available for trade. The allocation space is $X=[\underline{x},\overline{x}]$, where $\underline{x}>0$ denotes the minimal technologically feasible allocation. We interpret $x$ as a contractual allocation, while realized consumption $x^{\prime}$ is chosen by the buyer and is not directly contractible.
@@
This formulation implies that although the buyer purchases an allocation $x$, she optimally selects an actual consumption level $x^{\prime}\in[0,x]$. Any excess allocation $x-x^{\prime}$ is discarded without cost. Under zero marginal cost along the allocation dimension, the seller’s profit equals $p$.
@@
Assumption 3 requires that the virtual surplus at the efficient consumption level be nonnegative for every type $\theta$. Note that this assumption does not require the virtual surplus to be maximized [^4] And it cannot be the case. at the efficient consumption level, so the seller still has incentives to distort allocations. This assumption is imposed for equilibrium existence in the continuum‑type, continuous‑action baseline and does not drive the results; Appendix B shows the main constructions in a discretized action space without Assumption 3.
@@
This sufficient condition is satisfied if the lowest-type buyer’s efficient surplus is sufficiently large or if the buyer type distribution is sufficiently compressed.
@@ 2.2 Timing and Solution Concept
Acceptance terminates the game. If the buyer accepts in period $t$ ($a_{t}=1$), the buyer’s payoff is
@@
This constrained benchmark anticipates the efficiency constraint that emerges endogenously for the lowest type in our equilibrium constructions. Adapting the analysis from Lemma 3 yields the following characterization.
@@ 4.1 The Coasian Equilibrium
This particular outcome mirrors the standard Coasian benchmark in the following sense: every buyer type receives an efficient allocation with no delay, and the seller’s profit is driven to the surplus of the lowest buyer type. Since $u_{\min}<\pi(F)$, this payoff lies strictly below the static commitment benchmark.
We emphasize that, conditional on trade, the lowest buyer type $\underline{\theta}$ must receive an allocation $x\geq x^{e}(\underline{\theta})$ by the seller’s sequential rationality. However, it is not true that delay vanishes asymptotically as $\delta\to 1$ in all equilibria. A counterexample is presented in Section B.3.
@@ 4.3 The Folk Theorem
If $\underline{x}\geq x^{e}(\underline{\theta})$, then $\pi^{e}(F)=\pi(F)$ by Corollary 2 and $\pi^{e}(F)$ is the upper bound on all equilibrium payoffs. We therefore obtain the following corollary.
@@
Suppose $\underline{x}\geq x^{e}(\underline{\theta})$.
$u_{\min},\ \pi^{e}(F)]\ \supseteq\ SE(\delta)\quad\text{for all }\delta.$
The $\underline{x}\geq x^{e}(\underline{\theta})$ case is comparable to the “no-gap” case of Ausubel and Deneckere (1989b), where the lowest type has zero valuation. In both settings, the lowest type is indifferent across all feasible allocation outcomes, which in the setting of Ausubel and Deneckere (1989b) corresponds to the discounted probability of trade. We discuss whether $\pi^{e}(F)$ constitutes the upper bound on all equilibrium payoffs in more general settings in Appendix B.
@@ 4.4 Comparison
In the framework of Ausubel and Deneckere (1989b), where the lowest buyer valuation is normalized to zero (the “no-gap” case), the equilibrium payoff set extends to the full commitment payoff. Our analysis provides a complementary perspective on this result. In the setting of Ausubel and Deneckere (1989b), the only allocation dimension is the discounted probability of trade. When the lowest type has valuation zero in the no-gap case, the efficient timing of trade is set-valued for this type, including infinite delay. Thus, under limited commitment, once the lowest type is allocated infinite delay, the allocations for other types may be arbitrary, provided they are monotone in type. In this context, the upper bound of the Folk Theorem can be interpreted as the maximum profit achievable subject to the constraint that the lowest-type buyer receives an efficient time of trade (infinite delay). In their setting, this constrained maximum coincides with the unconstrained static commitment payoff.
Robustness to Markov Refinements.
We show that the Folk Theorem is robust to a class of Markov refinements that typically eliminate equilibrium multiplicity and restore the Coase conjecture in standard settings. This distinguishes our environment from those benchmarks. We begin by recalling the definition of weak-Markov equilibria.
@@
We now introduce a relaxation of this refinement.
@@
A common structural feature of these environments is that allocations are fixed, leaving price as the only screening instrument. In such settings, the on-path Markov refinement eliminates the seller’s ability to delay because the incentive to preempt future sales remains strong. In contrast, our framework allows the seller to choose both price and allocation, and the equilibrium path features decreasing allocations. This additional margin alters the incentives: accelerating sales requires offering the lower future allocation immediately, which reduces total surplus and therefore the seller’s profit. Hence, when agents are sufficiently patient, the incentive to maintain the equilibrium allocation sequence outweighs any gain from acceleration. The additional allocation margin therefore alters the force of Markov refinements, which restore selection in canonical settings but not in ours.
iscussion
5.1 Examples.
This section illustrates how the abstract primitives of the model arise in familiar economic environments. In the examples below, the seller chooses a one-dimensional allocation $x$, while the buyer selects realized utilization $x^{\prime}\leq x$ after purchase. Preferences take the form
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\bigl[v(x^{\prime})+\theta x^{\prime}\bigr],$
where $v(\cdot)$ is strictly concave.
While we view information goods as a particularly useful benchmark, the same allocation structure arises when the seller chooses access, scope, or quality with zero marginal cost, and realized usage is chosen ex post by the buyer with free disposability.
Consumer Data.
Consider an environment in which the allocation is a marketing list of consumer profiles ordered by decreasing valuation of a product. The buyer is a producer who can perfectly price discriminate across consumers once their profiles are accessed and who may acquire a larger list but optimally use only a sublist. If consumer profiles are indexed by $z$ and have valuations $v^{\prime}(z)$, then an allocation $x$ corresponds to access to all profiles $z\in[0,x]$. Realized utilization takes the form $x^{\prime}\leq x$, yielding profit
$u(x,\theta)=\int_{0}^{x}\bigl(v^{\prime}(z)-c\bigr)^{+}\,dz=\max_{0\leq x^{\prime}\leq x}\bigl[v(x^{\prime})+\theta x^{\prime}\bigr],$
where $c=-\theta$ is a constant production cost. The producer optimally serves only consumers whose valuations exceed the production cost. In this environment, the seller is a data broker who can change the allocation level $x$ at zero marginal cost after consumer profiles have been collected and aggregated. Although each type has a unique efficient utilization level, any allocation exceeding that level is allocatively efficient, which generates a set-valued efficient allocation correspondence.
Data-driven Decision Making.
Consider a buyer (e.g., an investor) facing an irreversible decision under uncertainty about $\omega\in\{-1,1\}$, each occurring with prior probability $1/2$. A type-$\theta$ buyer obtains payoff $\theta a\omega$ from choosing an action $a\in\{-1,1\}$ and aims to match $\omega$. Without data, expected utility is zero. After initial data collection, the seller offers datasets indexed by $x\in[0,1/2]$, interpreted as precision.
A dataset of precision $x$ enables the buyer to conduct a symmetric, costly experiment that refines the prior $(1/2,1/2)$ into posteriors (1/2x, 1/2+x)($^{\prime},\,1/2+x^{\prime}$) or (1/2+x, 1/2x)($^{\prime},\,1/2-x^{\prime}$), with $x^{\prime}\leq x$, each realized with probability $1/2$. Running the experiment incurs a cost $c(x^{\prime})$, where $c(\cdot)$ is strictly convex.
The buyer’s utility is
$u(x,\theta)=\max_{0\leq x^{\prime}\leq x}\left[2\theta x^{\prime}-c(x^{\prime})\right]=\max_{0\leq x^{\prime}\leq x}\left[v(x^{\prime})+\theta x^{\prime}\right],$
where $v(\cdot)$ is strictly concave.
In this example, free disposability arises because the buyer may optimally choose to run a less informative experiment than the dataset permits. Zero marginal cost reflects that, once created, supplying higher-precision datasets entails no additional cost. The efficient allocation is again set-valued.
5.2 When Efficient Allocations Are Unique.
@@
In this section, we analyze environments in which allocative efficiency is unique. We show that once efficiency becomes single-valued, the payoff multiplicity highlighted by the Folk Theorem disappears. To establish this point sharply, we adopt a discrete-type framework. As discussed above, the skimming property generally fails in the presence of quality differentiation, which makes a direct characterization of equilibrium strategies difficult. Addressing the problem in full generality is therefore beyond the scope of this paper.
@@
We first show that the equilibrium multiplicity highlighted in Theorem 2 depends on free disposability. In particular, when free disposability is absent, the equilibrium set collapses to a single point.
@@
In Proposition 1, the unique equilibrium payoff converges to $\pi^{e}(F)$ in the patient limit. This occurs because, in the two-type case, the efficient outcome coincides with the static monopoly outcome subject to the constraint that the low type receives the efficient allocation. In this setting, the low type $\theta_{2}$ receives the efficient allocation. Moreover, since $\theta_{1}$ is the high type, its allocation is not subject to downward distortion in static monopoly outcome, and it therefore also receives the efficient level.
On the other hand, when free disposability is present, the equilibrium payoff set contains $u_{\text{min}},\pi^{e}(F)]$. Thus, we have shown that equilibrium multiplicity depends on free disposability. Furthermore, attaining $\pi^{e}(F)$ is a special case driven by the two-type structure. In Appendix B.4, we show that $\pi^{e}(F)$ cannot be achieved in general without free disposability.
Positive Marginal Cost.
@@
Taken together, these results show that the payoff indeterminacy in our setting is driven by set-valued allocative efficiency, rather than by quality differentiation per se.
5.3 Menu Offers
Corollary 5.
By contrast, our baseline model rules out intratemporal discrimination. In this case, the maximal market-clearing profit is $u_{\min}$, which is strictly below the static commitment benchmark. This difference explains why the equilibrium outcomes and underlying economic forces in the present model differ from those in the existing literature.